You are asked by the owner of a local hotel chain to develop a customer satisfaction survey to determine the percentage of customers who are dissatisfied with service. In the past year, 20,000 customers were serviced. The owner desires a 95 percent level of confidence with an allowable statistical error of ±0.01. From past estimates, she believes that about 3.5 percent of customers have expressed dissatisfaction. What sample size should you use for this survey? Explain your answer. After the calculation, provide a brief explanation of which of the five key dimensions of service quality -- reliability, assurance, tangibles, empathy, or responsiveness -- should be emphasized in the satisfaction survey. Why?
Answer #1

Z value for 95 percent confidence level, Zα/2 =NORMSINV(1-(1-0.95)/2) = 1.96

Statistical error, E = 0.01

Fraction of customer expressed dissatisfaction, p̅ = 0.035

Sample size = z2*p̅*(1-p̅)/E2

= 1.962*0.035*(1-0.035)/0.012

= 1297

Sample size of 1297 should be used for this survey.